# On Neumaier Cayley graphs

### v1

@@ 1. Introduction
For every even positive integer $k$, there exists a strictly Neumaier Cayley graph with parameters
$(2^{2k+2},(2^{k+1}-1)(2^{k}+1),2(2^{k}-1)(2^{k-1}+1),2^{k},2^{k+1})$
over any abelian group of order $2^{2k+2}$ having an elementary abelian subgroup of index $2$.
@@ 4.2. Construction 22
Let $A_{0}\cong H_{0}\cong C_{2}^{k}$ and hence $G_{0}\cong C_{2}^{2k}$, and $l=2^{k-1}+1$. It is well-known that there exist pairwise disjoint and disjoint with $A_{0}$ subgroups $A_{1},\ldots,A_{l}$ of $G_{0}$ of order $2^{k}$. Put
@@
$(v_{0},k_{0},\lambda_{0},\mu_{0})=(2^{2k},(2^{k}-1)(2^{k-1}+1),2(2^{k-1}-1)(2^{k-2}+1),2^{k-1}(2^{k-1}+1))$
(see, e.g. [6]). Clearly, $S_{0}\cap A_{0}=\varnothing$. Due to the definition of $\Gamma_{0}$, each of the subgroups $A_{1},\ldots,A_{l}$ is a $2^{k-1}$-regular clique of size $2^{k}$ in $\Gamma_{0}$. So $\Gamma_{0}$ is a Neumaier graph.
Let $A$ be an abelian group of order $2^{k+1}$ such that $A\geq A_{0}$ and $|A:A_{0}|=2$, $C\cong C_{2}$, $H=H_{0}\times C\cong C_{2}^{k+1}$, and $G=A\times H$. By Lemma 3.1 (Lines 3 and 5 in Table 1), the group $H$ has an RSRDS $T$ with parameters $(2^{k},2,2^{k},2^{k-1})$ and forbidden subgroup $C$. Therefore one can construct the graph $\Gamma=\operatorname{Cay}(G,S)$ with connection set $S=(A\setminus A_{0})T\cup S_{0}C\cup\{c\}$. It is easy to verify that the parameters of $\Gamma_{0}$ satisfy Eq. (4) and hence $\Gamma$ is a Neumaier graph with the required parameters by Corollary 4.6. By Lemma 4.7, $\Gamma$ is a strictly Neumaier graph. ∎
@@
Algorithm 1 was realized using GAP and MAGMA. Applying Algorithm 1, a standard procedure of isomorphism testing between two graphs implemented in MAGMA, and counting the number of common neighbours of two non-adjacent vertices in each graph, we enumerate all pairwise nonisomorphic Neumaier Cayley graphs with a spread given by the cosets of a subgroup on at most 64 vertices and separate them into two parts consisting of strongly regular graphs and strictly Neumaier graphs. It turns out that most of the computed graphs have nexus 1 and there are only few Neumaier graphs with nexus greater than 1. All of the latter graphs are collected in Table 3. The smallest graph from Theorem 1.2 with parameters $(64,35,18,4,8)$ was found by a computer calculation using Algorithm 1.
Tuple #SRGs #SNGs
(16,9,4,2,4) - 1
(64,21,8,2,8) 11 -
(64,28,12,3,8) 119 18
(64,35,18,4,8) 726 1380
(64,42,26,5,8) 34 1
(64,49,36,6,8) 1 -
Table 3. Neumaier Cayley graphs with a spread given by the cosets of a subgroup and nexus >1>1 on at most 64 vertices

### v2

@@ 1. Introduction
For every even positive integer $n$, there exists a strictly Neumaier Cayley graph with parameters
$(2^{2n+2},(2^{n+1}-1)(2^{n}+1),2(2^{n}-1)(2^{n-1}+1),2^{n},2^{n+1})$
over any abelian group of order $2^{2n+2}$ having an elementary abelian subgroup of index $2$.
@@ 4.2. Construction 22
Let $n$ be a positive even integer, $A_{0}\cong H_{0}\cong C_{2}^{n}$ and hence $G_{0}\cong C_{2}^{2n}$, and $l=2^{n-1}+1$. It is well-known that there exist pairwise trivially intersecting and trivially intersecting with $A_{0}$ subgroups $A_{1},\ldots,A_{l}$ of $G_{0}$ of order $2^{n}$. Put
@@
$(v_{0},k_{0},\lambda_{0},\mu_{0})=(2^{2n},(2^{n}-1)(2^{n-1}+1),2(2^{n-1}-1)(2^{n-2}+1),2^{n-1}(2^{n-1}+1))$
(see, e.g. [6]). Clearly, $S_{0}\cap A_{0}=\varnothing$. Due to the definition of $\Gamma_{0}$, each of the subgroups $A_{1},\ldots,A_{l}$ is a $2^{n-1}$-regular clique of size $2^{n}$ in $\Gamma_{0}$. So $\Gamma_{0}$ is a Neumaier graph.
Let $A$ be an abelian group of order $2^{n+1}$ such that $A\geq A_{0}$ and $|A:A_{0}|=2$, $C\cong C_{2}$, $H=H_{0}\times C\cong C_{2}^{n+1}$, and $G=A\times H$. By Lemma 3.1 (Lines 3 and 5 in Table 1), the group $H$ has an RSRDS $T$ with parameters $(2^{n},2,2^{n},2^{n-1})$ and forbidden subgroup $C$. Therefore one can construct the graph $\Gamma=\operatorname{Cay}(G,S)$ with connection set $S=(A\setminus A_{0})T\cup S_{0}C\cup\{c\}$. It is easy to verify that the parameters of $\Gamma_{0}$ satisfy Eq. (4) and hence $\Gamma$ is a Neumaier graph with the required parameters by Corollary 4.6. By Lemma 4.7, $\Gamma$ is a strictly Neumaier graph. ∎
@@
Algorithm 1 was realized using GAP and MAGMA. Applying Algorithm 1, a standard procedure of isomorphism testing between two graphs implemented in MAGMA, and counting the number of common neighbours of two non-adjacent vertices in each graph, we enumerate all pairwise nonisomorphic strictly Neumaier Cayley graphs with a spread given by the cosets of a subgroup on at most 64 vertices. It turns out that most of the computed strictly Neumaier graphs have nexus 1 and there are only few strictly Neumaier graphs with nexus greater than 1. All of the latter graphs are collected in Table 3. The smallest graph from Theorem 1.2 with parameters $(64,35,18,4,8)$ was found by a computer calculation using Algorithm 1.
Tuple #SNGs
(16,9,4,2,4) 1
(64,28,12,3,8) 18
(64,35,18,4,8) 1380
(64,42,26,5,8) 1
Table 3. Strictly Neumaier Cayley graphs with a spread given by the cosets of a subgroup and nexus >1>1 on at most 64 vertices